Tom has taken 5 of the 8 equally weighted tests in his US History class. He has an average score of 78.0 points. How many points does he need to earn on the 6th test to bring his average score to 80 points?
Answer by ewatrrr(10682)

 
Hi
Tom has taken 5 of the 8 equally weighted tests in his US History class.
He has an average score of 78.0 points. Total point thru 5 tests = 5*78 = 390
How many points does he need to earn on the 6th test to bring his
average score to 80 points? Let x represent his points on his 6th test
(390 + x)/6 = 80
390 + x = 480
x = 90 points